Simplify these expressions and find their values if x = 3, a = – 1, b = – 2
 2 – 8x + 4x + 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that x = 3
We have,
= 2 + 4–8x + 4x
= 6–4x
Then, substitute the value of x in the equation
= 6 - (4 × 3)
= 6 - 12
=-6

Thanks